OpenStudy (lilmike234):

solution to the system of equations x+5y=13 y=2x-4

OpenStudy (tifferslovesyaoi):

x+5y=13 x+(2x-4)=13 x+2x-4=13 3x=17 x=5.666....... or x=5.67 I hope this is correct

OpenStudy (lilmike234):

So whats y=?